Physics 152, Spring 2019 Express your answer using three significant figures. gi 73
Answer #1

Answer : F3 on 2 = 18.0 N

91 42 73 2d

If r is the dtc between the 2 charges, force of attraction is

F = kq1q2/r^2

F1-2 = kq1q2/r12^2

F3-2 = kq3q2/r32^2

q3 = 8q1

and r32 = 2r12

from the above we get F3-2 = 18 N